# Jacobian-Enhanced Neural Networks

### v1

@@ Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization. This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
@@ I Introduction
Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output. The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and quantifies its benefits for the intended use-case. It complements previous work in the following ways. First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Finally, through judicious use of hyperparameters, it enables weighting; i.e. a process by which individual sample points can be prioritized differently in situations that call for it. The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve SBO in gradient-descent applications.
@@ I-A Intended Use-Case
JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. This opens the door to useful design applications, such as Monte Carlo simulations of computationally expensive functions for example. However, this is true for any type of surrogate model in general. In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: Surrogate-Based Optimization (SBO). The field of aerospace engineering is rich in design applications of such a use-case, with entire book chapters dedicated to it [1, 2] and applied examples commonly found in aerodynamic applications [3, 4, 5].
@@ I-C Supporting Software
An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as tensorflow [7], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [8]. However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, and requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap; it is intended for those engineers in a rush with a need to predict partials accurately and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [9] is JENN with a different API (present author contribution).
@@ II Previous Work
A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient Python libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [9], and even review papers about gradient-enhanced meta-models do not mention neural networks, except in passing [10]. However, early work in the late 1990’s can be found [11] which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. Those authors conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
@@
Follow-on research by that group provides the mathematical derivation of the equations for neural net gradient-enhancement [12], but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome in the late 2000’s by a deep, neural-network formulation with gradient-enhancement [3], in which the authors apply their algorithm to surrogate-based optimization of airfoil designs using evolutionary algorithms. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
@@
Recent work in 2019 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools [4]. Those authors successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not counter-productive in the first place. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions. Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
@@
Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [13] (PINN). This creative research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator). However, PINN is outside the target scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
@@ III-D Parameter Update
where the hyper-parameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such ADAM [15], but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
@@ III-E Backward Propagation
Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.r.t.) the inputs . By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
III-E1 Loss Function Derivatives w.r.t. Neural Net Parameters
@@
Substituting these expressions into Eq. 11 yields:
III-E2 Loss Function Derivatives w.r.t. to Neural Net Activations
@@
The derivatives for the next layer are obtained through the chain rule. According to Eq. III-D, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, current layer derivatives w.r.t. to previous layer activations are:
@@
All necessary equations to compute the derivatives of the loss function w.r.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
III-E3 Cost Function Derivatives w.r.t. Parameters
@@ IV Verification and Validation
The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier. This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig. 2, which compares JENN against a standard Neural Net (NN) when training data is sparse (shown as dots). Higher dimensional examples are available on the project’s homepage, but were not included since it is hard to visualize more than two dimensions on a static plot. All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data (provided the cost of obtaining the Jacobian is not counter-productive). Finally, it was also verified that runtime scales as using the Rastrigin function to generate progressively larger datasets and training a new model, as shown in Fig. 2c using a last generation MacBook Pro (2021).
@@ V Example Application: Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural net, to optimize some slow-running function of interest, rather than directly using the function itself. This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[16]. The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization. This will be accomplished by solving the following optimization problem:
@@
where is a surrogate-model approximation to the famous Rosenbrock function:
Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3 where optimizer convergence histories are shown for the true response and its surrogate-model approximations using JENN and NN under the same settings, except for jacobian-enhancement, respectively. Training data (not shown to avoid clutter) was generate by sampling the true response on a regular grid. Training histories are provided in Figs. 3a–3b to verify that both neural networks converged to the same degree.
@@ V-A Results Without Polishing
In a first step, consider Fig.3c which does not use polishing (explained in the next section). Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed. For both surrogates, the terminal value is indeed close to zero (the true optimum) but neither JENN nor NN reach the true location of this optimum situated at ; although JENN does get much closer. The Rosenbrock function is notorious for exhibiting a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization. Carefully studying the contour lines, one can indeed observe subtle undulations in this shallow valley within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
This finding is not surprising. Recall from Eq. III-D that neural nets learn by minimizing prediction error. Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting small features in the vicinity of the optimum, where prediction error implies negigible penalty. This situation is at odds with the needs of gradient-based optimization, which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
@@ V-B Results With Polishing
For most practical engineering problems, close is good enough. However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design (to ensure smooth contours free of undulations). This situation can be ameliorated by using what will be called polishing. It consists of restarting a previously trained model and continue training it after magnifying some portion of the design space to be improved. This is accomplished by judiciously setting or to prioritize specific training points belonging to that subspace. For example, a radial basis function centered on zero could be used to magnify small numbers in the aforementioned shallow valley of the Rosenbrock function:
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are , whereas nears those points . This has the desired effect of disproportionately emphasizing small values of and during training by up to a factor of 1000. The results are shown in Fig. 3d where it can be seen that JENN now recovers the true optimum almost exactly; artificial local minima nearly vanish. This is not true for NN; although results improve noticeably. Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
@@ VI Noisy Partials
The answer to this question is shown by example in Fig. 4 using the Rastrigin function as a test bed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural net (gradient-enhancement off). Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
@@ VII Conclusion
In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions. It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place. It was further shown that JENN outperforms standard neural nets in the context of surrogate-based optimization, getting much closer to the true optimum but not quite reaching it for the selected example. In fact, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero. However, it was shown that this shortcoming can be overcome through a process termed polishing, which prioritizes salient sample points in that neighborhood through judicious use of hyperparameters. This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it. The open-source Python library associated with this work was provided as a link in earlier sections, in the hope that it will be useful to others and spark new contributions in this researcher area.

### v2

@@ Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models. Since a surrogate emulates the original model accurately in real time, it yields a computational speed benefit that can be used to carry out orders of magnitude more function calls quickly. However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization. This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
@@ I Introduction
Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output. The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and quantifies its benefits for the intended use-case. It complements previous work in the following ways. First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Finally, through judicious use of hyperparameters, it enables weighting; i.e. a process by which individual sample points can be prioritized differently in situations that call for it. The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve Surrogate-Based Optimization (SBO) in gradient-descent applications.
@@ I-A Intended Use-Case
JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models. Since a surrogate emulates the original model accurately in fractions of second, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. This allows designers to explore a broader tradespace, and more readily quantify the impact of various parameters in order to better understand the performance of the system under design. While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives with improved accuracy. This is a critical property for one important use-case: surrogate-based optimization. The field of engineering is rich in design applications of such a use-case with entire book chapters dedicated to it [1, 2, 3, 4], review papers [5, 6, 7, 8], and countless examples to be found, particularly in the field of aerospace engineering [9, 10, 11, 12].
@@ I-C Supporting Software
An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as TensorFlow [14] and PyTorch [15], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [16]. However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap. It is intended for engineers with limited time, who have a need to predict partials accurately, and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [17] is JENN with a different API (present author contribution).
@@ II Previous Work
A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [17], and even review papers about gradient-enhanced meta-models only mention neural networks in passing [18]. However, early work by Sellar and Batill [19] can be found in the late 1990’s, which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. They conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
@@
Follow-on research by that Liu and Batill [20] provides the mathematical derivation of the equations for neural net gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome by Giannakoglou et al. [10] in the late 2000’s who developed a deep neural-network formulation with gradient-enhancement and use it for aerodynamic design applications. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
@@
Recent work by Bouhlel et al. [11] in 2020 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools. Bouhlel et al. successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not prohibitive in the first place. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions. Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
The works surveyed so far have focused on prediction accuracy, but Hoffman et al. [21] take a different approach by focusing on robustness instead. They introduce the concept of Jacobian regularization for classification, where the key idea is to ensure robustness of models against input perturbations due to adversarial attacks; for example, a purposefully modified input image to avoid detection. However, robustness is outside the scope of the present work, which is to predict design partials accurately given high quality inputs.
@@
Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [22] (PINN). This novel research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator). However, PINN is outside the scope of the present work, which is to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
@@ III-D Parameter Update
where the hyper-parameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such as ADAM [24], but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least rquares Estimator (LSE), augmented with an extra term to account for partial derivatives:
@@ III-E Backward Propagation
Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.s.t.) the inputs . By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
III-E1 Loss Function Derivatives w.s.t. Neural Net Parameters
@@
rubstituting these expressions into Eq. 11 yields:
III-E2 Loss Function Derivatives w.s.t. to Neural Net Activations
@@
The derivatives for the next layer are obtained through the chain rule. According to Eq. III-D, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, current layer derivatives w.s.t. to previous layer activations are:
@@
All necessary equations to compute the derivatives of the loss function w.s.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
III-E3 Cost Function Derivatives w.s.t. Parameters
@@ IV Verification and Validation
The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier. This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig. 2, where training data is shown as dots. These validation results compare JENN against a standard Neural Net (NN), defined as the equivalent neural network without gradient-enhancement.
All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data, so long as the computational cost of obtaining the Jacobian is not prohibitive in the first place. In Figs. 2a and 2b, where there are only three and four training samples respectively, NN fails to even capture the correct trend whereas JENN predicts the response almost perfectly. In Fig. 2d, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction. By contrast, under the same conditions, JENN predicts the Rastrigin function [25] almost perfectly.
Finally, it was also verified that runtime scales as , as shown in Fig. 2c. This plot was obtained using the Rastrigin function to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
@@ V Example Application: Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural networks, to optimize some slow-running function of interest, rather than directly using the function itself. This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[26]. The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization. This will be accomplished by solving the following optimization problem:
@@
where is a surrogate-model approximation to the Rosenbrock [27] function, which is used as an optimization benchmark because it exhibits a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization:
Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3b, where optimization histories are shown for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished). Training data is not shown to avoid clutter, but was generated by sampling the true response on a regular grid plus 100 latin hypercube points. The neural network training histories are also included in Fig. 3a, in order to verify that both neural networks converged to the same degree.
@@ V-A Results Without Polishing
Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed: neither JENN nor NN reach the true location of this optimum situated at . Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer. These effects are less pronounced for JENN, but the optimizer gets trapped nonetheless.
In hindsight, this finding is not surprising because neural networks learn by minimizing prediction error (recall Eq. III-D). Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting flat regions in the vicinity of a zero optimum, where prediction error penalties are negligible. This situation is at odds with the needs of gradient-based optimization which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate. However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design where it is imperative that the final airfoil is free of undulations and non-smooth contours. An imperative research question therefore arises: how can these competing needs be reconciled?
@@ V-B Results With Polishing
This situation can be ameliorated through “polishing.” Given a trained model, this approach consists of train it further after magnifying regions of interest. This is accomplished by judiciously setting to prioritize specific training points belonging to that subspace. For example, flat regions could be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on such as:
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are whereas near those points . This has the desired effect of disproportionately emphasizing regions of small slope during training by up to a factor of 1000 at the expense of de-emphasizing other regions, which is why it is necessary to start from an already trained model.
Final results are shown in Fig. 3b for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly. Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
@@ VI Noisy Partials
The answer to this question is shown by example in Fig. 4 using the Rastrigin function as a testbed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural network (gradient-enhancement off). Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
@@ VII Conclusion
In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions. It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place. It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization. However, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero. It was further shown that this shortcoming can be overcome through polishing, which can be used to prioritize salient regions of interest through judicious use of hyperparameters. This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it. The ideas and findings of this paper have been published as an open-source library to allow readers to validate the results, and hopefully assist them in their surrogate modeling efforts.